OpenStudy (samsan9):

The equation for a projectile's height versus time is h(t)=-16t^2+vot+ho. A tennis ball machine serves a ball vertically into the air from a height of 2 feet, with an initial speed of 130 ft per second. Which equation correctly models the ball's height as a function of time?

OpenStudy (samsan9):

A.) h(t)=-16t^2-130t+2 B.) h(t)= -16t^2+130t+2 C.) h(t)= -16t^2+130 D.) h(t)= -16t^2-2t+130